OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (2, 7) and(-2, 9)

OpenStudy (jackellyn):

y=mx+b Solve for the slope, m y2-y1 9-7 2 1 m=---- = ---- = ---- = --- x2-x1 -2-2 -4 -2 7=(-1/2)(2) + b 7 = -1 + b 8 = b y= (-1/2)x + 8